# Mathematical Language Models: A Survey

### v3

@@ 4.2. Tool-based Methods
To improve the performance of mathematical reasoning, math-specific tools such as symbolic solvers and programs are utilized for LLMs (He-Yueya et al., 2023; Chen et al., 2022a). For example, SymbLLM (He-Yueya et al., 2023) solved math word problems by combining language models with symbolic solvers. They focus on automatically generating high-quality, step-by-step solutions to mathematical word problems, especially those encountered in mathematical applications.
Furthermore, previous studies have explored the process of converting mathematical problems into code (Chen et al., 2022a; Drori et al., 2022). PoT (Chen et al., 2022a) proposes a fusion of CoT with programs, while Drori et al. (Drori et al., 2022) showcased the effectiveness of a neural network pre-trained on text and fine-tuned on code, particularly leveraging OpenAI’s Codex transformer. This approach successfully solves, explains and generates math problems at the university level. MathPrompter (Imani et al., 2023) employs a zero-shot chain-of-thought prompting technique to generate multiple algebraic expressions or Python functions in varied ways to solve the same math problem, enhancing reliance on output results. PAL (Gao et al., 2023) introduces an innovative approach to bolster the performance of pre-trained language models (PLMs) in mathematical problem-solving. This involves utilizing LLMs to comprehend natural language problems and generate programs as intermediate reasoning steps.
In addition to solving math problems, programs can also play a role in tutoring math. For instance, Upadhyay explores the role of large language models in tutoring systems (Upadhyay et al., 2023), introducing a “code scratchpad” alongside the traditional “language scratchpad” to enhance the model’s performance in tutoring steps, particularly using a grade school mathematics dataset.
@@
To facilitate the seamless integration of LLMs with diverse tools, several multi-tool approaches have been proposed to enable LLMs to learn how to use multiple tools simultaneously. Toolformer (Schick et al., 2023) adopts a self-supervised training approach to enable the utilization of different tools such as search engines, calculators and translation systems via simple API calls. This is accomplished by fine-tuning on a vast collection of sampled API calls and filtering based on their ability to reduce perplexity on subsequent tokens. Chameleon (Lu et al., 2023a), on the other hand, enhances LLMs with plug-and-play modules designed for compositional reasoning. It creates programs by combining various tools, including LLMs, off-the-shelf vision models, web search engines, Python functions, and heuristic-based modules, to tackle complex reasoning tasks. The Automatic Reasoning and Tool-use (ART) framework (Paranjape et al., 2023) leverages frozen LLMs to automatically generate intermediate reasoning steps as a program. This seamlessly incorporates external tools to support computations that surpass the core capabilities of LLMs. Meanwhile, ToolkenGPT (Hao et al., 2023b) adopts a strategy of learning “toolken” embeddings to represent each tool as a token. This empowers LLMs to effortlessly utilize tools similar to generating word tokens. It caters to a broader range of tools and utilizes extensive demonstration data to learn toolkit embeddings.
Furthermore, CRITIC (Gou et al., 2023a) significantly enhances the outputs of LLMs by enabling them to verify and self-correct through interactions with external tools. Inspired by human cognition and critical thinking, CRITIC continuously refines text generated by LLMs through iterative interactions with tools like search engines and code interpreters. Tool Augmented Language Models (TALM) (Parisi et al., 2022) seamlessly integrate language models with non-differentiable tools. They employ an iterative “self-play” method to bootstrap performance based on a limited number of tool demonstrations. In contrast, Tool-Documentation (Hsieh et al., 2023) opted for tools based on documentation rather than relying on demonstrations. Current practices involve teaching LLMs through a few-shot demonstration of tool usage, a process prone to bias and overfitting. The alternative proposed by Tool-Documentation, utilizing tool documentation that provides descriptions of individual tool usage, is argued to be a more effective approach. ToRA (Gou et al., 2023b) employs a set of Tool-integrated Reasoning Agents, which integrate computational libraries and symbolic solvers, among other mathematical tools, to effectively solve complex mathematical problems.
@@ 4.3. Fundamental CoT Methods
Foundation of CoT.
In the initial stages of research, a limited body of work has leveraged the principles of chain-of-thought (CoT) to enhance the mathematical capabilities of language models. Notably, Ling et al (Ling et al., 2017) propose a methodology involving the generation of a series of concise steps, termed ”Answer Rationales,” to guide the resolution of algebraic word problems. MathQA (Amini et al., 2019) suggests decomposing Math Word Problems into multiple steps corresponding to programmatic operations for resolution.
Utilizing the in-context learning of LLMs, CoT (Wei et al., 2022) explicitly introduces the concept of the chain of thought (CoT) for the first time, and substantiates its efficacy in enhancing reasoning abilities. A CoT example is shown in Figure 8, where the text highlighted in blue and green guides to the ultimately correct answer. This validation is also corroborated in Complexity-based CoT (Fu et al., 2023). Following a similar line of thought, LVSS (Lightman et al., 2023) presents a Process Reward Model (PRM) and compares it with the Outcome Reward Model (ORM). Empirical findings affirm that process supervision outperforms outcome supervision, leading to notable performance improvements on the MATH (Hendrycks et al., 2021b) benchmarks. Consequently, several methods (Chen et al., 2022a; Yuan et al., 2023a; Yu et al., 2023a; Wang et al., 2023a; Sel et al., 2023; Yue et al., 2023) leverage CoT methodologies to enhance the mathematical performance of models.
Figure 8. Example of CoT in Wei et al.(Wei et al., 2022).
@@
RFT (Yuan et al., 2023a) proposes the application of rejection sampling finetuning (RFT) to gather more optimal reasoning paths, thereby enhancing mathematical reasoning performance. Thought Propagation (Yu et al., 2023a) considers gaining insights from analogous problems to assist in addressing the current issue. Wang et al. (Wang et al., 2023a) proposed a Fine-Tuning (AFT) paradigm that aligns the model to prioritize the generation of CoT with superior performance. Additionally, PoT (Chen et al., 2022a) introduces a Program of Thoughts, combining CoT and programming, and AOT (Sel et al., 2023) enhances reasoning abilities through algorithmic-style demonstrations of CoT. Furthermore, MAmmoTH (Yue et al., 2023) integrates CoT and PoT (Chen et al., 2022a) rationales to instruct large language models in utilizing code tools for solving mathematical problems.
Construction of CoT.
To streamline the process of CoT creation, various approaches (Lu et al., 2023b; Zhang et al., 2023c; Fu et al., 2023; Wang et al., 2023a; Kojima et al., [n. d.]; Shum et al., 2023; Pitis et al., 2023; Sel et al., 2023; Yu et al., 2023a) have been introduced for its automatic generation.
@@
Zero-shot CoT (Kojima et al., [n. d.]) introduces a simple CoT example, “Let’s think step by step,” which effectively enhances the model’s reasoning capabilities. Auto-CoT (Zhang et al., 2023c) selects representative questions through clustering and answers them using “Let’s think step by step,” concatenating the question-answer pairs one by one to automatically generate CoT. In the case of Complexity-based CoT (Fu et al., 2023), the demonstration is chosen by simply selecting the reasoning chain with the highest number of inference steps sampled from the model. Using reinforcement learning, PromptPG-CoT (Lu et al., 2023b) trains a model with policy gradient to assist GPT-3 in selecting suitable CoT demonstrations. Similarly, AutoMate CoT (Shum et al., 2023) employs a variance-reduced policy gradient strategy to estimate the significance of each example in a black box language model, thereby selecting more effective examples. BoostedPrompt (Pitis et al., 2023) presents an iterative prompt ensemble method that enhances prompts when the current demonstration faces challenges in handling specific problems.
@@ 4.4. Advanced CoT Methods
LLMs often produce incorrect reasoning steps, which can lead to a series of cascading errors in their solutions. Implementing verification and refining the reasoning process based on feedback can significantly reduce these errors. This approach is akin to human reflection and involves critically evaluating each step of the reasoning process. Various verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023) have been proposed to address these issues.
Zhou et al. (Zhou et al., 2023d) propose a code-based self-verification approach, utilizing a zero-shot prompt to encourage the GPT-4 Code Interpreter to employ code for self-verifying its responses, enhancing its mathematical reasoning capabilities. To mitigate the challenge of validating the entire deductive reasoning process, VerifyCoT (Ling et al., 2023) introduces a deductive reasoning form, ensuring that each reasoning step strictly relies on the preceding steps. Furthermore, DIVERSE (Li et al., 2023b) independently verifies each reasoning step and a voting mechanism to eliminate incorrect answers. Both Verify-and-Edit (Zhao et al., 2023a) and Retrieval-CoT (He et al., 2022) utilizes external retrieval tools to support the model in validating reasoning rationales. The key difference is the former edits rationales using retrieved information, while the latter helps the model self-rethink to improve performance on complex reasoning tasks. Both methods effectively reduce factual mistakes during the reasoning process. Shridhar et al. (Shridhar et al., 2023) summarize the use of the revisions approach in the SCREWS framework, which includes a selection module to choose between the original and modified reasoning steps.
@@
Due to the inherent stochastic nature of language models, which output probability distributions over predicted words, they may randomly generate incorrect reasoning steps and outcomes. To tackle this challenge, some methods (Wang et al., 2023b; Naik et al., 2023; Fu et al., 2023; Li et al., 2023b; Miao et al., 2023; Yoran et al., 2023; Cobbe et al., 2021; Khalifa et al., 2023) leverage the concept of ensemble learning, employing techniques such as voting and ranking to eliminate uncertainties in the reasoning process.
Self-Consistency (Wang et al., 2023b) employs multiple reasoning paths and selects the final response through a simple majority vote. Similarly, Diversity-of-Thought (Naik et al., 2023) generates diverse reasoning paths by altering prompts and aggregates the responses via the majority vote. Complexity-based CoT(Fu et al., 2023) favors answers derived from more intricate reasoning paths. DIVERSE (Li et al., 2023b) uses a weighted voting mechanism to filter out incorrect answers. Nevertheless, these voting-based methods often overlook the potentially useful information within unsuccessful CoT reasoning and lack an efficient integration of multiple reasoning chains to improve performance. Self-check (Miao et al., 2023) addresses this by incorporating reasoning steps into the voting mechanism, ensuring both consistent answers and reliable reasoning. MCR (Yoran et al., 2023) takes a step further by consolidating information across various reasoning chains, acquiring the most relevant facts, facilitating a more comprehensive analysis to make successful reasoning.
Additionally, there are a few methods based on ranking. Rank-verifier (Cobbe et al., 2021) proposes using a ranking system to judge the correctness of model completions. Furthermore, GRACE (Khalifa et al., 2023) leverages a discriminator trained via contrastive learning to rank each reasoning step.
@@
The original structure of the Chain of Thought (CoT) is sequential, facing limitations when handling highly complex problems and lacking the ability for retrospective correction. To make the CoT chain structure more systematic and intricate, certain planning-based methods (Yao et al., 2023b; Mo and Xin, 2023; Long, 2023; Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023; Jiang et al., 2023a; Hao et al., 2023a; Zhou et al., 2023e; Liu et al., 2023a; Dagan et al., 2023; Madaan et al., 2023; Sun et al., 2023; Zhou et al., 2023b; Shinn et al., 2023) have been proposed. These methods achieve this by altering the organizational structure of reasoning steps or incorporating mechanisms for refinement and reflection.
@@
The Tree-of-Thought (ToT) (Yao et al., 2023b) proposes organizing reasoning paths into a tree structure, where each node represents a reasoning step, and edges denote dependencies between nodes. During the inference process, ToT can use self-assessment to determine future actions. This structure facilitates both forward and backward exploration, employing either deep-first or breadth-first search techniques. TouT (Mo and Xin, 2023) effectively employs Monte Carlo Dropout to assess uncertainty scores associated with diverse local responses of language models at intermediate steps, enhancing response precision through integration with global search algorithms. Long et al. (Long, 2023) introduced a multi-module ToT system, where the state of the ToT thought chain process can be stored and retraced through its storage module. Furthermore, Yao et al. (Yao et al., 2023a), Besta et al. (Besta et al., 2023), and Lei et al. (Lei et al., 2023) introduced a novel thought structure called the Graph of Thought (GoT). Similar to ToT, each graph node represents a distinct thought step or state, with edges indicating dependencies between nodes. These methods can effectively integrate diverse reasoning thoughts, fostering collaborative outcomes and improving reasoning by incorporating feedback loops. This contributes to an overall enhancement in the inference capacity of the thought network. Additionally, RESPROMPT (Jiang et al., 2023a) introduces Residual Connection Prompting, which converts the input prompt into complex reasoning graphs.
Inspired by Monte Carlo tree search, RAP (Hao et al., 2023a) utilizes the Monte Carlo tree search algorithm to traverse tree structures during the reasoning process. This approach effectively balances global and local searches to obtain a high-reward reasoning path. On the other hand, LATS (Zhou et al., 2023e) employs LLMs as agents, value functions, and optimizers, repurposing their inherent strengths to enhance decision-making. Liu et al. (Liu et al., 2023a) and Dagan et al. (Dagan et al., 2023) introduced LLM+P and LLM+DP, incorporating the generation of Planning Domain Definition Language (PDDL) to break down complex problems and execute plans using specialized models.
Furthermore, the self-refine approach, as proposed by Madaan et al. (Madaan et al., 2023) and Sun et al. (Sun et al., 2023), focuses on error correction and summarizing past experiences. Specifically, Self-Refine performs pre-execution validation, providing feedback to the model for refinement in case of errors. Zhou et al. (Zhou et al., 2023b) then took a step further by merging the iterative self-refine with PDDL in ISR-LLM. Additionally, Shinn et al. (Shinn et al., 2023) introduced re-flexion to rectify errors stemming from previous actions.
@@
The Socratic teaching method (Nelson, 1980), derived from the philosophies of ancient Greek thinker Socrates, provides a unique perspective when considering the integration of LMs in mathematics education (Davies et al., 2021). This method, emphasizing questioning and dialogue, aims to guide learners toward profound thinking (Davis and Steinglass, 1997; Brickhouse and Smith, 2009). In realm of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery.

### v4

@@ 4.2. Tool-based Methods
To improve the performance of mathematical reasoning, math-specific tools such as symbolic solvers and programs are utilized for LLMs (He-Yueya et al., 2023; Chen et al., 2022a). For example, SymbLLM (He-Yueya et al., 2023) solved MWPs by combining LMs with symbolic solvers. They focus on automatically generating high-quality, step-by-step solutions to mathematical word problems, especially those encountered in mathematical applications. LeanReasoner (Jiang et al., 2024) and Lean Copilot (Song et al., 2024) integrates LLMs with the Lean theorem proving framework. Leveraging Lean’s symbolic solver reduces the risk of logical inconsistencies and enhances the LLMs’ capability to handle complex reasoning tasks through the extensive theorem libraries available in Lean. Similarly, LINC (Olausson et al., 2023) integrates LLMs with a first-order logic theorem prover. The LLM translates natural language premises and conclusions into First-Order Logic expressions. These expressions are then sent to an external theorem prover, such as Prover9, for symbolic deductive reasoning to determine the truth value of the conclusion. Furthermore, previous studies have explored the process of converting MWPs into code (Chen et al., 2022a; Drori et al., 2022). PoT (Chen et al., 2022a) proposes a fusion of CoT with programs, while Drori et al. (Drori et al., 2022) showcased the effectiveness of pre-training on text and fine-tuning on code. This approach successfully solves, explains and generates math problems at the university level. MathPrompter (Imani et al., 2023) employs a zero-shot CoT prompting technique to generate multiple algebraic expressions or Python functions in varied ways to solve the same math problem. PAL (Gao et al., 2023) introduces an innovative approach to bolster the performance of LLMs in mathematical problem-solving. This involves utilizing LLMs to comprehend natural language problems and generate programs as intermediate reasoning steps. To enhance the ability of LLMs to solve MWPs using code, MuMath-Code (Yin et al., 2024) constructed an augmented dataset that uses multi-perspective data augmentation methods and synthesizes code-nested solutions. In addition to solving math problems, programs can also play a role in tutoring math. For instance, Upadhyay explores the role of LLMs in tutoring systems (Upadhyay et al., 2023), introducing a “code scratchpad” alongside the traditional “language scratchpad” to enhance the model’s performance in tutoring steps, particularly using a grade school mathematics dataset.
@@
To facilitate the seamless integration of LLMs with diverse tools, several multi-tool approaches have been proposed to enable LLMs to learn how to use multiple tools simultaneously. Toolformer (Schick et al., 2023) adopts a self-supervised training approach to enable the utilization of different tools such as search engines, calculators and translation systems via simple API calls. This is accomplished by fine-tuning on a vast collection of sampled API calls and filtering based on their ability to reduce perplexity on subsequent tokens. Chameleon (Lu et al., 2023b), on the other hand, enhances LLMs with plug-and-play modules designed for compositional reasoning. It creates programs by combining various tools, including LLMs, off-the-shelf vision models, web search engines, Python functions, and heuristic-based modules, to tackle complex reasoning tasks. The Automatic Reasoning and Tool-use (ART) framework (Paranjape et al., 2023) leverages frozen LLMs to automatically generate intermediate reasoning steps as a program. This seamlessly incorporates external tools to support computations that surpass the core capabilities of LLMs. Meanwhile, ToolkenGPT (Hao et al., 2023b) adopts a strategy of learning “toolken” embeddings to represent each tool as a token. This empowers LLMs to effortlessly utilize tools similar to generating word tokens. It caters to a broader range of tools and utilizes extensive demonstration data to learn toolkit embeddings. Furthermore, CRITIC (Gou et al., 2023a) significantly enhances the outputs of LLMs by enabling them to verify and self-correct through interactions with external tools. Inspired by human cognition and critical thinking, CRITIC continuously refines text generated by LLMs through iterative interactions with tools like search engines and code interpreters. Tool Augmented Language Models (TALM) (Parisi et al., 2022) seamlessly integrate language models with non-differentiable tools. They employ an iterative “self-play” method to bootstrap performance based on a limited number of tool demonstrations. In contrast, Tool-Documentation (Hsieh et al., 2023) opted for tools based on documentation rather than relying on demonstrations. Current practices involve teaching LLMs through a few-shot demonstration of tool usage, a process prone to bias and overfitting. The alternative proposed by Tool-Documentation, utilizing tool documentation that provides descriptions of individual tool usage, is argued to be a more effective approach. ToRA (Gou et al., 2023b) employs a set of Tool-integrated Reasoning Agents, which integrate computational libraries and symbolic solvers, among other mathematical tools, to effectively solve complex MWPs.
@@ 4.3. Fundamental CoT Methods
Foundation of CoT
@@
In the initial stages of research, a limited body of work has leveraged the principles of CoT to enhance the mathematical capabilities of language models. Notably, Ling et al (Ling et al., 2017) propose a methodology involving the generation of a series of concise steps, termed ”Answer Rationales,” to guide the resolution of algebraic word problems. MathQA (Amini et al., 2019) suggests decomposing Math Word Problems into multiple steps corresponding to programmatic operations for resolution. Utilizing the in-context learning of LLMs, CoT (Wei et al., 2022) explicitly introduces the concept of the CoT for the first time, and substantiates its efficacy in enhancing reasoning abilities. This validation is also corroborated in Complexity-based CoT (Fu et al., 2023). Consequently, several methods (Chen et al., 2022a; Yuan et al., 2023a; Yu et al., 2023a; Wang et al., 2023a; Sel et al., 2023; Yue et al., 2023) leverage CoT methodologies to enhance the mathematical performance of models. RFT (Yuan et al., 2023a) proposes the application of rejection sampling finetuning (RFT) to gather more optimal reasoning paths, thereby enhancing mathematical reasoning performance. Thought Propagation (Yu et al., 2023a) considers gaining insights from analogous problems to assist in addressing the current issue. Wang et al. (Wang et al., 2023a) proposed a Fine-Tuning (AFT) paradigm that aligns the model to prioritize the generation of CoT with superior performance. SymbCoT (Xu et al., 2024a) integrates symbolic expressions and logical rules with CoT prompting, enhancing the logical rigor of CoT. Additionally, PoT (Chen et al., 2022a) introduces a Program of Thoughts, combining CoT and programming, and AOT (Sel et al., 2023) enhances reasoning abilities through algorithmic-style demonstrations of CoT. Furthermore, MAmmoTH (Yue et al., 2023) integrates CoT and PoT (Chen et al., 2022a) rationales to instruct large language models in utilizing code tools for solving mathematical problems.
Construction of CoT
@@
To streamline the process of CoT creation, various approaches (Lu et al., 2023c; Zhang et al., 2023e; Fu et al., 2023; Wang et al., 2023a; Kojima et al., [n. d.]; Shum et al., 2023; Pitis et al., 2023; Sel et al., 2023; Yu et al., 2023a) have been introduced for its automatic generation. Zero-shot CoT (Kojima et al., [n. d.]) introduces a simple CoT example, “Let’s think step by step,” which effectively enhances the model’s reasoning capabilities. Auto-CoT (Zhang et al., 2023e) selects representative questions through clustering and answers them using “Let’s think step by step,” concatenating the question-answer pairs one by one to automatically generate CoT. In the case of Complexity-based CoT (Fu et al., 2023), the demonstration is chosen by simply selecting the reasoning chain with the highest number of inference steps sampled from the model. Several methods aim to enhance the quality of CoT generated by models. CoT-Influx (Huang et al., 2024) trains a pruner specifically optimized for enhancing the mathematical capabilities of CoT generation, and validates the effectiveness of the pruner in improving the mathematical performance of LLMs on the GSM8K (Cobbe et al., 2021). Using reinforcement learning, PromptPG-CoT (Lu et al., 2023c) trains a model with policy gradient to assist GPT-3 in selecting suitable CoT demonstrations. Similarly, AutoMate CoT (Shum et al., 2023) employs a variance-reduced policy gradient strategy to estimate the significance of each example in a black box language model, thereby selecting more effective examples. BoostedPrompt (Pitis et al., 2023) presents an iterative prompt ensemble method that enhances prompts when the current demonstration faces challenges in handling specific problems.
@@ 4.4. Advanced CoT Methods
LLMs often produce incorrect reasoning steps, which can lead to a series of cascading errors in their solutions. Implementing verification and refining the reasoning process based on feedback can significantly reduce these errors. This approach is akin to human reflection and involves critically evaluating each step of the reasoning process. Various verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023) have been proposed to address these issues. Zhou et al. (Zhou et al., 2023d) propose a code-based self-verification approach, utilizing a zero-shot prompt to encourage the GPT-4 Code Interpreter to employ code for self-verifying its responses, enhancing its mathematical reasoning capabilities. To mitigate the challenge of validating the entire deductive reasoning process, VerifyCoT (Ling et al., 2023) introduces a deductive reasoning form, ensuring that each reasoning step strictly relies on the preceding steps. Furthermore, DIVERSE (Li et al., 2023b) independently verifies each reasoning step and a voting mechanism to eliminate incorrect answers. Both Verify-and-Edit (Zhao et al., 2023a) and Retrieval-CoT (He et al., 2022) utilizes external retrieval tools to support the model in validating reasoning rationales. The key difference is the former edits rationales using retrieved information, while the latter helps the model self-rethink to improve performance on complex reasoning tasks. Both methods effectively reduce factual mistakes during the reasoning process. Shridhar et al. (Shridhar et al., 2023) summarize the use of the revisions approach in the SCREWS framework, which includes a selection module to choose between the original and modified reasoning steps. SSC-CoT (Zhao et al., 2024a) found that prompting the model from critical intermediate steps can enhance the success rate of solving mathematical problems. Therefore, it emphasizes verifying these key steps during the inference process.
@@
Due to the inherent stochastic nature of LLMs, which output probability distributions over predicted words, they may randomly generate incorrect reasoning steps and outcomes. To tackle this challenge, some methods (Wang et al., 2023d; Naik et al., 2023; Fu et al., 2023; Li et al., 2023b; Miao et al., 2023; Yoran et al., 2023; Cobbe et al., 2021; Khalifa et al., 2023) leverage the concept of ensemble learning, employing techniques such as voting and ranking to eliminate uncertainties in the reasoning process. Self-Consistency (Wang et al., 2023d) employs multiple reasoning paths and selects the final response through a simple majority vote. Similarly, Diversity-of-Thought (Naik et al., 2023) generates diverse reasoning paths by altering prompts and aggregates the responses via the majority vote. Complexity-based CoT(Fu et al., 2023) favors answers derived from more intricate reasoning paths. DIVERSE (Li et al., 2023b) uses a weighted voting mechanism to filter out incorrect answers. Nevertheless, these voting-based methods often overlook the potentially useful information within unsuccessful CoT reasoning and lack an efficient integration of multiple reasoning chains to improve performance. Self-check (Miao et al., 2023) addresses this by incorporating reasoning steps into the voting mechanism, ensuring both consistent answers and reliable reasoning. MCR (Yoran et al., 2023) takes a step further by consolidating information across various reasoning chains, acquiring the most relevant facts, facilitating a more comprehensive analysis to make successful reasoning. Additionally, there are a few methods based on ranking. Rank-verifier (Cobbe et al., 2021) proposes using a ranking system to judge the correctness of model completions. Furthermore, GRACE (Khalifa et al., 2023) leverages a discriminator trained via contrastive learning to rank each reasoning step.
@@
The original structure of the Chain of Thought (CoT) is sequential, facing limitations when handling highly complex problems and lacking the ability for retrospective correction. To make the CoT chain structure more systematic and intricate, certain planning-based methods (Yao et al., 2023b; Mo and Xin, 2023; Long, 2023; Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023; Jiang et al., 2023a; Hao et al., 2023a; Zhou et al., 2023e; Liu et al., 2023a; Dagan et al., 2023; Madaan et al., 2023; Sun et al., 2023; Zhou et al., 2023b; Shinn et al., 2023) have been proposed. These methods achieve this by altering the organizational structure of reasoning steps or incorporating mechanisms for refinement and reflection. The Tree-of-Thought (ToT) (Yao et al., 2023b) proposes organizing reasoning paths into a tree structure, where each node represents a reasoning step, and edges denote dependencies between nodes. During the inference process, ToT can use self-assessment to determine future actions. This structure facilitates both forward and backward exploration, employing either deep-first or breadth-first search techniques. TouT (Mo and Xin, 2023) effectively employs Monte Carlo Dropout to assess uncertainty scores associated with diverse local responses of language models at intermediate steps, enhancing response precision through integration with global search algorithms. Long et al. (Long, 2023) introduced a multi-module ToT system, where the state of the ToT thought chain process can be stored and retraced through its storage module. Furthermore, Yao et al. (Yao et al., 2023a), Besta et al. (Besta et al., 2023), and Lei et al. (Lei et al., 2023) introduced a novel thought structure called the Graph of Thought (GoT). Similar to ToT, each graph node represents a distinct thought step or state, with edges indicating dependencies between nodes. These methods can effectively integrate diverse reasoning thoughts, fostering collaborative outcomes and improving reasoning by incorporating feedback loops. This contributes to an overall enhancement in the inference capacity of the thought network. CR (Zhang et al., 2023c) proposes a method where LLMs assume three roles—Proposer, Verifier, and Reporter—to collaboratively solve problems. This approach differs from CoT and ToT in its ability to dynamically store and utilize all historically validated inference results for execution and combination, a concept akin to GoT. Additionally, RESPROMPT (Jiang et al., 2023a) introduces Residual Connection Prompting, which converts the input prompt into complex reasoning graphs. Inspired by Monte Carlo tree search (MCTS) (Chaslot et al., 2008), RAP (Hao et al., 2023a) utilizes the MCTS to traverse tree structures during the reasoning process. This approach effectively balances global and local searches to obtain a high-reward reasoning path. On the other hand, LATS (Zhou et al., 2023e) employs LLMs as agents, value functions, and optimizers, repurposing their inherent strengths to enhance decision-making. Step-Plan (Zhang et al., 2023b) propose a step-by-step planning approach for intermediate solution generation, which strategically plans the generation of the next solution step. Liu et al. (Liu et al., 2023a) and Dagan et al. (Dagan et al., 2023) introduced LLM+P and LLM+DP, incorporating the generation of Planning Domain Definition Language (PDDL) to break down complex problems and execute plans using specialized models. Furthermore, Self-Refine, as proposed by Madaan et al. (Madaan et al., 2023) and Sun et al. (Sun et al., 2023), focuses on error correction and summarizing past experiences. Specifically, Self-Refine performs pre-execution validation, providing feedback to the model for refinement in case of errors. Zhou et al. (Zhou et al., 2023b) then took a step further by merging the iterative self-refine with PDDL in ISR-LLM. Additionally, Shinn et al. (Shinn et al., 2023) introduced re-flexion to rectify errors stemming from previous actions.
Self-correct Methods.
Recently, a lot of Self-correct Methods have been proposed, which introduce a self-correction mechanism on top of CoT. Combined with techniques like MCTS and RL, these methods enable the model to iteratively enhance itself. The success of OpenAI’s o1 (OpenAI, 2024) marks a significant advancement in high-level reasoning by employing long CoT and self-correction mechanisms. Building on this, QwQ (Team, 2024), Marco-o1 (Zhao et al., 2024c) and o3 have also been introduced, further validating the effectiveness of long CoT and self-correction. STaR (Zelikman et al., 2022) leverages existing LLM reasoning to iteratively create a high-quality dataset, refining the model through self-generated inferences. Similar to STaR, REFT (Trung et al., 2024) does not employ SFT, and instead refines the model using the PPO in an online reinforcement learning setting. V-STaR (Hosseini et al., 2024) builds upon STaR by retaining error trajectories during the iterative generation process. Quiet-STaR (Zelikman et al., 2024) enables LLMs to generate a “rationale” before predicting the next token, explaining the anticipated content of the future text. CoRe (Zhu et al., 2022) mimics the human brain’s dual-system model, separating reasoning into generation and verification, and it uses MCTS for feedback during inference. Additionally, SCoRe (Kumar et al., 2024) introduces a two-stage reinforcement strategy to mitigate distribution shifts and behavior collapse, facilitating self-correction. MCTSr (Zhang et al., 2024b) proposes a method based on Monte Carlo Tree Self-Refine that enhances LLM’s performance on the Mathematical Olympiad, matching GPT-4. Flow-DPO (Deng and Mineiro, 2024) introduces a method where two LLM agents collaborate to generate step-level preference data, which is used to train the LLM using the DPO algorithm. Lightman et al. (Lightman et al., 2024) presents a Process Reward Model (PRM) and compares it with the Outcome Reward Model. Step-DPO (Lai et al., 2024) also introduces a method in which each step of the reasoning process is optimized based on preferences, rather than evaluating the entire answer as a whole. OmegaPRM (Luo et al., 2024) proposes an automated process supervision method with a divide-and-conquer style MCTS algorithm, which significantly improves Gemini’s proficiency in mathematical reasoning. Moreover, HGS-PRM (Ma et al., 2023) proposes combining Heuristic Greedy Search with PRM to optimize the reasoning paths.